• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 17:28:16 on May 14, 2008

    Entry number 237980

    keyword=Status of Hall A on Wednesday 14 May 2008

    The status of Hall A is Restricted Access.

    The target work has been finished and the scattering chamber is being
    pumped down. J. P. Chen has been contacted about the cooldown tomorrow.
    I have not heard back from him.

    The crane was fixed and the detector packages were swapped with the
    "electron" package now on BigBite. BitBite is being recabled.

    The plan for the night is to finish up cabling BigBite and take cosmics
    overnight with BigBite.

    To Do:

    1) Make sure the polarity of the magnets/spectrometers are set correctly.
    2) Leak check & cool down target.
    3) hacsbc2 is behaving strangely. What's up with that!?
    4) Lock-up hall
    5) Post check-out plan once feedback has been received from someone/anyone.